Calculate the osmotic pressure at 273k of a 5%solution of urea.(mole mass of urea=60)

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
Since 5% solution of urea means that amount of urea, NH2CONH2, w =5g and Volume of solution,V =100ml  By relation, Osmotic pressure, P=wRT/MV
P=5*0.0821*273 / 60*100
P=112.6*1000 / 6000
P=18.68atm